The total current in a parallel circuit equals the sum of the currents in the branches. The total current in six parallel branches is 1.9A. Some of the branches have currents of 0.25A and others have 0.45A. How many of each type of branch are in the circuit?
Answer by stanbon(75887)

The total current in a parallel circuit equals the sum of the currents in the branches. The total current in six parallel branches is 1.9A. Some of the branches have currents of 0.25A and others have 0.45A. How many of each type of branch are in the circuit?
----
Equations:
Quantity: t + f = 6
Amps::::0.25t + 0.45f = 1.9
---
Multiply thru Quantity by 25
Multiply thru Amps by 100
----
25t + 25f = 25*6
25t + 45f = 190
----
Subtract and solve for "f":
20f = 40
f = 2 (# of 0.45A circuits)
---
Solve for "t":
t + f = 6
t = 4 (# of 0.25A circuits)
===============================
Cheers,
Stan H.